As the chief financial officer of Blades, Inc., Ben Holt is pleased that his current system of exporting "Speedos" to Thailand seems to be working well. Blades' primary customer in Thailand, a retailer called Entertainment Products, has committed itself to purchasing a fixed number of Speedos annually for the next 3 years at a fixed price denominated in baht, Thailand's currency. Furthermore, Blades is using a Thai supplier for some of the components needed to manufacture Speedos. Nevertheless, Holt is concerned about recent developments in Asia. Foreign investors from various countries had invested heavily in Thailand to take advantage of the high interest rates there. As a result of the weak economy in Thailand, however, many foreign investors have lost confidence in Thailand and have withdrawn their funds. Holt has two major concerns regarding these developments. First, he is wondering how these changes in Thailand's economy could affect the value of the Thai baht and, consequently, Blades. More specifically, he is wondering whether the effects on the Thai baht may affect Blades even though its primary Thai customer is committed to Blades over the next 3 years. Second, Holt believes that Blades may be able to speculate on the anticipated movement of the baht, but he is uncertain about the procedure needed to accomplish this. To facilitate Holt's understanding of exchange rate speculation, he has asked you, Blades' financial analyst, to provide him with detailed illustrations of two scenarios. In the first, the baht would move from a current level of $.022 to $.020 within the next 30 days. Under the second scenario, the baht would move from its current level to $.025 within the next 30 days. How are percentage changes in a currency's value measured? Illustrate your answer numerically by assuming a change in the Thai baht's value from a value of $.022 to $.026. What are the basic factors that determine the value of a currency? In equilibrium, what is the relationship between these factors?
Answer #1

Firstly, factors influencing value of a currency relative to another (in this case, value of the Thai Baht relative to USD) are as follows:

1. Interest rates in both the countries, that is, interest rates in USA and Thailand.

2. Inflation rates in both the countries

3. Governments performance in Thailand, including Central bank's intervention in the forex market and foreign currency debt taken by the government

4. Speculation activities on Thai Baht

As per the Covered Interest Rate parity, which assumes market to be in equilibrium, exchange rate of a currency relative to another currency is a factor of the interest rates in both the countries. Higher the interest rate, more is the exchange rate expected to depreciate.

Scenario 1: if exchange rate moves from $0.022 to $0.02 per Baht

Value of the USD relative to the Baht have improved, in other words, the Thai Baht has depreciated against the USD.

$0.022 per Baht in terms of USD is Baht 45.45 per USD

$0.02 per Baht in terms of USD is Baht 50 per USD

Depreciation = 45.45 - 50 = -4.55 per USD

Depreciation (%) = -4.55/45.45 = 10%

Scenario 2: if exchange rate moves from $0.022 to $0.025 per Baht

Value of the USD relative to the Baht have fallen, in other words, the Thai Baht has appreciated against the USD.

$0.022 per Baht in terms of USD is Baht 45.45 per USD

$0.025 per Baht in terms of USD is Baht 40 per USD

Appreciation = 45.45 - 40 = 5.45 per USD

Appreciation (%) = 5.45/45.45 = 12%

Note: Due to Thailand's weak economy and loosing investor confidence, there could be a chance of investors withdrawing their investments from Thailand which would hit Thailand's currency negatively due to foreign currency outflows from the country.